“Gasohol” is a mixture of gasoline and ethyl alcohol used as automobile fuel. The alcohol produces energy in a combustion reaction with O2.
C2H5OH(l) + 3O2(g) → 2CO2(g) + 3H2O(l)
If 0.115 g of alcohol evolves 3.42 kJ when burned at constant pressure, what is the molar enthalpy of combustion for ethyl alcohol?

Chemical cold packs contain two pouches, one containing solid ammonium nitrate and the other containing water (often dyed for no real reason). When the contents of the pouches mix, the ammonium nitrate dissolves in the water, NH4NO3(s) → NH4+(aq) + NO3-(aq)
The dissolving process is endothermic. The reaction absorbs approximately 326 J/g of ammonium nitrate. To lower the temperature of 220. g of water by 20.0°C, how many grams of ammonium nitrate would you need to dissolve?
1
Expert's answer
2015-11-02T12:13:59-0500
Molar mass of ethanol is 46 g / mol. Then in the combustion of 1 mole of ethanol is highlighted
46 * 3.42 / 0.115 = 1368 kJ

The heat capacity of water is 4.18 J / g ° K
To lower the temperature of 220 grams of water at 20.0 ° C, to absorb 4.18 * 220 * 20 = 18392 J.
To this must be dissolved in water
18392/326 = 56.4 grams of ammonium nitrate